Exercise 6 (if not specific mentioned, answers should be in 4 decimals or simplified; highlight your final answer by color 1. just need 1 and 3, please use matlab codes
Answer #1

Given finna sin ( 22 ) WCOS 727) 21-) (29-31 fm= cos ) fm=cos () (C13) (2x 39, *) (22) fm= cos () (**pen to y felay = 11 (3. (a) Given fini= 423+5(1-1.5² flui= 4(342) +5X2 (-15) (1-0) flim = 1242 +10 (1-15) fthl=12n +100-15 For critical points t